a student has 2 small books, five medium sized books, and three large ones. If all the books are different, how many ways can they be arranged on a shelf so that all books of the same size are together.
Answer by checkley77(12844)

The 2 small books have 2*1=2 arrangements.
The 5 medium books have 5*4*3*2*1=120 arrangements.
The 3 latge books fave 3*2*1=6 arrangemenys.
3*2*1=6 different ways to arrange these 3 groups of books.
Thus the total=2*120*6*6=8,640 different arrangements.